OpenStudy (anonymous):

Suppose the population of a town is 8,600 and is growing 3% each year. Predict the population after 3 years. A. about 77,400 people B. about 9,397 people C. about 26574 people D. about 232,200 people

OpenStudy (anonymous):

use compound interest formulae:)

OpenStudy (tkhunny):

Only B is even remotely sensible. The other three are immediately ridiculous.

OpenStudy (anonymous):

8600+3%=8600+(3*8600/100)=8600+258=8858(first year) 8858+3%=8858+(3*8858/100)=8858+265.74=9123.74 9123.74+3%=9123.74+(3*9123.74/100)=9123.74+273.7122=9397.4522 so 9397 peoples:)

OpenStudy (tkhunny):

Ever hear of the Rule of 72? It's a funny thing with remarkably broad applciation. It is this, divide 72 but the %rate of increase and you get approximately the number of years to double. In this case, we have 3%, so doubling is about 72/3 = 24 years. Thus, we should be nowhere near 17000 by year 3. This eliminates all but 9397. Exact Answer, 8600*(1.03)^3 = 9397.4522, but most folks don't like 0.4522 people, so round to integer.